(4/5) ÷ 4 is equal to                                   
 (4/5) ÷ 4 is equal to                                   
 (4/5) ÷ 4 is equal to                                   

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(4/5) ÷ 4 is equal to 1/5.
= (4/5) ÷ 4

= (4/5) × (1⁄4)
= (1/5) × (1/1)
= 1/5



Thanks